# Deep Learning Calabi-Yau four folds with hybrid and recurrent neural network architectures

### v2

@@ Abstract
In this work, we report the results of applying deep learning based on hybrid convolutional-recurrent and purely recurrent neural network architectures to the dataset of almost one million complete intersection Calabi-Yau four-folds (CICY4) to machine-learn their four Hodge numbers $h^{1,1}$, $h^{2,1}$, $h^{3,1}$, $h^{2,2}$. In particular, we explored and experimented with twelve different neural network models, nine of which are convolutional-recurrent (CNN-RNN) hybrids with the RNN unit being either GRU (Gated Recurrent Unit) or Long Short Term Memory (LSTM). The remaining four models are purely recurrent neural networks based on LSTM. In terms of the $h^{1,1}$, $h^{2,1}$, $h^{3,1}$ and $h^{2,2}$ prediction accuracies, at 72% training ratio, our best performing individual model is CNN-LSTM-400, a hybrid CNN-LSTM with the LSTM hidden size of 400, which obtained 99.74%, 98.07%, 95.19%, 81.01%, our second best performing individual model is LSTM-448, an LSTM-based model with the hidden size of 448, which obtained 99.74%, 97.51%, 94.24%, and 78.63%. These results were improved by forming ensembles of the top two, three or even four models. Our best ensemble, consisting of the top four models, achieved the accuracies of 99.84%, 98.71%, 96.26%, 85.03%. At 80% training ratio, the top two performing models LSTM-448 and LSTM-424 are both LSTM-based with the hidden sizes of 448 and 424. Compared with the 72% training ratio, there is a significant improvement of accuracies, which reached 99.85%, 98.66%, 96.26%, 84.77% for the best individual model and 99.90%, 99.03%, 97.97%, 87.34% for the best ensemble.
@@ 1 Introduction
In this work, we move away from the aforementioned established approach of using purely convolutional architecture for deep learning the CICY Hodge numbers. Instead, our approach relies mainly on variants of recurrent neural networks (RNNs) such as Gated Recurrent Unit (GRU) [40], [41], [42] and Long-Short Term Memory (LSTM) [36], [37], [39], [38] in two different settings. The first setting involves hybrid architectures based on CNN and RNN (both GRU and LSTM). The second setting involves purely recurrent architecture based on LSTM. By experimenting with different variations in each of the two architectures above, we obtain a number of promising results that are comparable to those obtained in [25]. Ultimately, with these results, we hope to demonstrate and establish the effectiveness of recurrent neural networks in this particular problem which has been primarily explored with convolutional architectures.
Before moving on to the next section, we would like to quickly highlight some recent works that employ AI/ML in the study of different types of CY manifolds. The work [12] studies the fine, regular, star triangulations (FRSTs) of two reflexive polytopes (one with $h^{1,1}$ = 15, $h^{2,1}$ = 42, and the other with $h^{1,1}=491$, $h^{2,1}=11$) from the Kreuzer Skarke dataset [11] that contains around 473 million four-dimensional reflexive polytopes whose triangulations correspond to the ambient toric varieties in which CY three-folds can be defined as hypersurfaces by employing a neural network made of ReZero layers [32] to learn several geometric quantities. The work [16] applied both unsupervised and supervised machine learning methods to the database of weighted $\mathbb{P}^{4}$ admitting CY 3-fold hypersurfaces to study the CY topological data. The work [13] (see also [14]) used neural networks to compute the numerical Ricci-flat metrics for CICY and Kreuzer-Skarke CY manifolds. The work [15] uses genetic algorithms generates new reflexive polytopes in various dimensions used to construct the toric varieties hosting the CY manifolds as hypersurfaces. The work [17] uses ANNs to machine-learn the Hodge numbers of CY five-folds constructed as hypersurfaces in weighted projective spaces. The work [18] uses ANNs and symbolic regression to machine learn the topological quantities related to the Sasakian and $G_{2}$-geometries of contact Calabi-Yau 7-manifolds. The reviews [19], [20] cover an extensive list of earlier works on the topic of machine learning CY manifolds in various contexts.
@@
The organization of this note is as follows: In Section 2, we summarize the main characteristics of the dataset and our data preparation method. In section 3, we recall the basic mathematical facts about recurrent neural network architecture before moving to section 4 in which we describe in detail the different types of neural networks architectures used in this work. In section 5, we report the training results obtained for all 12 models at 72% data split (see sections 5.1, 5.2, 5.3). In section 5.4, we form four ensembles of the best performing models and check their performances. Next, the accuracies of all models (evaluated on the test set), including the ensembles, are presented in section 5.5. In section 6, we perform a 5-fold cross validation training for a single model and use the results (section 6.1) to form new ensembles (section 6.2). In section 7.1, we report the training results of the top three individual models which were retrained on the enlarged dataset at 80% data split. In section 7.2, we again form several new ensembles and check the performances of the retrained models and and their associated ensembles (see section 7.3). A summary in section 8 is followed by the Appendix 9 containing the sections 9.1, 9.2, 9.3 on the additional metrics for evaluating the performance of the neural networks, and the section 9.4 that includes the training curves of all models. The Python codes in the form of Jupyter notebooks  [^2] We make use of the following standard libraries: numpy (https://numpy.org/) [48], pandas (https://pandas.pydata.org/) [49], scikit-learn (https://scikit-learn.org) [50], matplotlib (https://matplotlib.org/) [52], seaborn (https://seaborn.pydata.org/) [53] for data handling, manipulation, visualization, and Pytorch (https://pytorch.org/) [51] for building and training the neural networks. for this work can be found at the following GitHub link: https://github.com/lorrespz/CICY4-Deep-learning-hybrid-recurrent-NNs-main.
@@ 2.1 Overview of the CICY4 dataset
The analysis of [28] excluded 15,813 CICY4 manifolds on the account of them being product manifolds, so the dataset actually contains 905,684 CICY4 manifolds and their corresponding Hodge numbers. The mean, maximal and minimal values for the Hodge numbers are reported in [28]
@@ 2.2 Data preparation
The originally compiled dataset by the authors of [24] can be downloaded in either text format or as a Mathematica file at [33]. A more convenient format suitable for deep learning using Python code is Numpy array npy, which was created by the authors of [25], [34]. The npy files can be obtained by running the script [34] on the original text file downloaded from [33]. The final dataset format used for the deep learning task thus consists of 905,684 data points of $16\times 20$ configuration matrices (whose entries are rescaled to be within the range $[0,1]$) and their associated Hodge number tuples ($h^{(1,1)}$,$h^{(2,1)}$, $h^{(3,1)}$, $h^{(2,2)}$). As noted in [25], there is a heavy imbalance in the dataset for $h^{2,1}$, as well as for $h^{3,1}$ and $h^{2,2}$: A large portion of the dataset for these Hodge numbers (on the order of $10^{5}$ data points) having values in the lower ranges. Furthermore, $h^{2,2}$ and $h^{3,1}$ (to a lesser extent, $h^{2,1}$) contain many outliers in the tail end of the data. These characteristics make the regression task for $h^{3,1}$, $h^{2,2}$ much harder than that for $h^{2,1}$ and $h^{1,1}$. Known results from the previous works [30], [25] have demonstrated the fact that although perfect accuracies could be obtained for $h^{1,1}$ and $h^{2,1}$, the accuracies for $h^{3,1}$ and $h^{2,2}$ are usually much lower.
@@
  • For the 5-fold CV training, the same datasets as the 72% data split was used, but with the following crucial differences:
    • The training set (72% of the data) is merged with the validation set (8% of the data), to form a total training set accounting for 80% of the data. This total training set is converted into 5 training datasets corresponding to the 5 CV training folds. Each ’fold’ (where ’fold’ now is used synonymously with the training set for that fold) contains five partitions. In each of the training fold, a model is trained on the data of the 4 partitions (accounting for 64% of the full data) and validated on the remaning partition (accounting for 16% of the full data). For a 5-fold CV, there are 5 models. Although these models all have the same architecture, they should be treated as independent models, since they are trained and validated on different subsets of the training data.
@@ 3 Basics of recurrent neural networks
In this section, we briefly recall some pertinent mathematical facts about general recurrent neural networks (RNNs) and their more sophisticated GRU/LSTM variants for the benefits of the physicist readers who might not be familiar with deep learning. While our work makes use of both convolutional and recurrent architectures, we will not touch on the basics of convolutional neural networks, since many excellent texts already exist on CNN on the topics of deep learning CY manifolds as CNNs are the main tool for deep learning in these works (see, for example, the works [1], [5], [6], [4] which also discuss the basics of deep learning using a simple feed forward/dense articial neural network in great detail).
@@ 5 Training results (using 72% dataset)
In this section, we report the results obtained after the first round of training in which all twelve models was trained on the 72% dataset. Our training is done entirely on Kaggle (an AI/ML competition hosting platform) - https://www.kaggle.com - using the P100 GPU unit, which comes with a time usage limit777of 12 hours per notebook runtime session, (see https://www.kaggle.com/docs/notebooks) as well as a maximum of 30 hours per week, renewable at the start of each week. See https://www.kaggle.com/docs/efficient-gpu-usage. All networks were trained using HuberLoss [^8] https://pytorch.org/docs/stable/generated/torch.nn.HuberLoss.html, AdamW optimizer [^9] https://pytorch.org/docs/stable/generated/torch.optim.AdamW.html and a batch size of 128. The use of HuberLoss for training ensures a robust regression performance in the presence of outliers since HuberLoss combines the advantages of both the Mean Squared Error loss and the Mean Absolute Error (MAE) loss by using a $\delta$ parameter as follows:
@@ 5.1 CNN-GRU hybrid neural networks
  • For comparison among all CNN-GRU models, see Fig.16 which shows the validation accuracies obtained during training for all models in both normal and logarithmic scales. The training curves showing both training and validation losses for individual CNN-GRU models can be found in Fig.38.
    Figure 16: Validation losses of all CNN-GRU models (in normal and logarithmic scale).
@@ 5.2 CNN-LSTM hybrid neural networks
  • The best performing CNN-LSTM hybrid model is CNN-LSTM-400 (with the train/test results noted in bold font). This is also the best model out of the 12 models considered in this work. A notable trend of these models’ results is the relatively stable performance observed at $M=256$ and $M=284$. Peak performance is attained at $M=400$. When increasing $M$ further to 416, the performance drops dramatically due to very slow convergence. In fact, the worst performing model of all models considered is CNN-LSTM-416, which also has the largest number of parameters ($3.017\times 10^{6}$). While all models achieved convergence after an average of around 300 epochs, CNN-LSTM-416’s training still progresses at epoch 550 (the cutoff). If training was allowed to continue, the model would presumably continue to improve, but due to the time constraint with Kaggle’s GPU usage, it was not possible to extend the training time.
  • The obviously superior performance of LSTM-based hybrid networks over GRU-based ones can be seen when comparing between CNN-GRU-384 and CNN-LSTM-384 (Fig.17).
    Figure 17: Validation accuracies of CNN-LSTM-384 and CNN-GRU-384 networks.
  • For comparison among the six CNN-LSTM models, see Fig.18 which shows the validation losses (normal and logarithmic scales) of all CNN-LSTM models. The training curves showing training and validation losses for each individual CNN-LSTM model can be found in Fig.39 (in Appendix 9.4).
    Figure 18: Validation losses of all CNN-LSTM models (normal and log scales). The lowest loss value (the green curve) corresponds to the best model CNN-LSTM-400. The red curve, corresponding to CNN-LSTM-416, did not drop to zero but ended at a very high value at the end of epoch 550 (signaling a failure to converge).
@@ 5.3 LSTM-based neural networks
  • For comparison among the four LSTM-based models, the validation losses of all four models are shown in Fig.19 in normal and logarithmic scales.
    Figure 19: Validation losses of all LSTM models (normal and log scales). In the log plot, the green curve with the smallest loss corresponds to LSTM-448.
  • The training curves for individual LSTM-based models showing training and validation losses can be found in Fig.40.
@@ 5.4 Ensembles of best models
Forming ensemble is a common technique to improve the overall performance, especially when there are several trained models ready to be deployed. Due to the fact that each model has extracted their own useful features and thus learnt a different representation of the data, they might have different performance strengths in different parts of the data distribution. For instance, a model that performs well on a certain range of the data might not do well on a different range of the same data. In this sense, the diversity of model architectures in the ensemble potentially leads to a more robust performance on the entire data set. In an ensemble, the predictions of all the constituent models are averaged, either evenly by taking the mean - as is the case in this work, or unevenly by taking an weighted average.
For the task at hand, the specific ensembles that we consider are listed in Table 5, together with their associated test accuracies (best results are listed in bold).

### v3

@@ Abstract
In this work, we report the results of applying deep learning based on hybrid convolutional-recurrent and purely recurrent neural network architectures to the dataset of almost one million complete intersection Calabi-Yau four-folds (CICY4) to machine-learn their four Hodge numbers $h^{1,1}$, $h^{2,1}$, $h^{3,1}$, $h^{2,2}$. In particular, we explored and experimented with twelve different neural network models, nine of which are convolutional-recurrent (CNN-RNN) hybrids with the RNN unit being either GRU (Gated Recurrent Unit) or Long Short Term Memory (LSTM). The remaining four models are purely recurrent neural networks based on LSTM. In terms of the $h^{1,1}$, $h^{2,1}$, $h^{3,1}$ and $h^{2,2}$ prediction accuracies, at 72% training ratio, our best performing individual model is CNN-LSTM-400, a hybrid CNN-LSTM with the LSTM hidden size of 400, which obtained 99.74%, 98.07%, 95.19%, 81.01%, our second best performing individual model is LSTM-448, an LSTM-based model with the hidden size of 448, which obtained 99.74%, 97.51%, 94.24%, and 78.63%. These results were improved by forming ensembles of the top two, three or even four models. Our best ensemble, consisting of the top four models, achieved the accuracies of 99.84%, 98.71%, 96.26%, 85.03%. At 80% training ratio, the top two performing models LSTM-448 and LSTM-424 are both LSTM-based with the hidden sizes of 448 and 424. Compared with the 72% training ratio, there is a significant improvement of accuracies, which reached 99.85%, 98.66%, 96.26%, 84.77% for the best individual model and 99.90%, 99.03%, 97.97%, 87.34% for the best ensemble. By nature a proof of concept, the results of this work conclusively established the utility of RNN-based architectures and demonstrated their effective performances compared to the well-explored purely CNN-based architectures in the problem of deep learning Calabi Yau manifolds.
@@ 1 Introduction
Moreover, the fact that many aspects of CY manifolds are amenable to AI/ML techniques has opened up many different avenues for new research in this direction. In particular, the recent work [12] studies the fine, regular, star triangulations (FRSTs) of two reflexive polytopes from the Kreuzer Skarke dataset [11] that contains around 473 million four-dimensional reflexive polytopes whose triangulations correspond to the ambient toric varieties in which CY three-folds can be defined as hypersurfaces by employing a neural network made of ReZero layers [32] to learn several geometric quantities. The work [16] applied both unsupervised and supervised machine learning methods to the database of weighted $\mathbb{P}^{4}$ admitting CY 3-fold hypersurfaces to study the CY topological data. The work [13] (see also [14]) used neural networks to compute the numerical Ricci-flat metrics for CICY and Kreuzer-Skarke CY manifolds. The work [15] uses genetic algorithms generates new reflexive polytopes in various dimensions used to construct the toric varieties hosting the CY manifolds as hypersurfaces. The work [17] uses ANNs to machine-learn the Hodge numbers of CY five-folds constructed as hypersurfaces in weighted projective spaces. The work [18] uses ANNs and symbolic regression to machine learn the topological quantities related to the Sasakian and $G_{2}$-geometries of contact Calabi-Yau 7-manifolds. The reviews [19], [20] cover an extensive list of earlier works on the topic of machine learning CY manifolds in various contexts.
@@
In this work, we move away from the aforementioned established approach of using purely convolutional architecture for deep learning the CICY Hodge numbers. Instead, our approach relies mainly on variants of recurrent neural networks (RNNs) such as Gated Recurrent Unit (GRU) [40], [41], [42] and Long-Short Term Memory (LSTM) [36], [37], [39], [38] in two different settings. The first setting involves hybrid architectures based on CNN and RNN (both GRU and LSTM). The second setting involves purely recurrent architecture based on LSTM. By experimenting with different variations in each of the two architectures above, we obtain a number of promising results that are comparable to or even outperforming those obtained in [25]. Ultimately, with these results, we hope to demonstrate and establish the effectiveness of recurrent neural networks in this particular problem which has been primarily and exclusively explored with convolutional architectures.
@@
The organization of this paper is as follows: In Section 2, we summarize the main characteristics of the dataset and our data preparation method. In section 3, we recall the basic mathematical facts about recurrent neural network architecture before moving to section 4 in which we describe in detail the different types of neural networks architectures used in this work. In section 5, we report the training results obtained for all 12 models at 72% data split (see sections 5.1, 5.2, 5.3). In section 5.4, we form four ensembles of the best performing models and check their performances. Next, the accuracies of all models (evaluated on the test set), including the ensembles, are presented in section 5.5. In section 6, we perform a 5-fold cross validation training for a single model and use the results (section 6.1) to form new ensembles (section 6.2). In section 7.1, we report the training results of the top three individual models which were retrained on the enlarged dataset at 80% data split. In section 7.2, we again form several new ensembles and check the performances of the retrained models and and their associated ensembles (see section 7.3). A summary in section 8 is followed by the Appendix 9 containing the sections 9.1, 9.2, 9.3 on the additional metrics for evaluating the performance of the neural networks, and the section 9.5 that includes the training curves of all models. The Python codes in the form of Kaggle notebooks  [^2] We make use of the following standard libraries: numpy (https://numpy.org/) [48], pandas (https://pandas.pydata.org/) [49], scikit-learn (https://scikit-learn.org) [50], matplotlib (https://matplotlib.org/) [52], seaborn (https://seaborn.pydata.org/) [53] for data handling, manipulation, visualization, and Pytorch (https://pytorch.org/) [51] for building and training the neural networks. for this work can be found at the following GitHub link: https://github.com/lorrespz/CICY4-Deep-learning-hybrid-recurrent-NNs-main.
@@ 2.1 Overview of the CICY4 dataset
The analysis of [28] and [25] excluded 15,813 CICY4 manifolds on the account of them being product manifolds, so the dataset actually contains 905,684 CICY4 manifolds and their corresponding Hodge numbers. The mean, maximal and minimal values for the Hodge numbers are reported in [28]
@@ 2.2 Data preparation
The originally compiled dataset by the authors of [24] can be downloaded in either text format or as a Mathematica file at [33]. A more convenient format suitable for deep learning using Python code is Numpy array npy, which was created by the authors of [25], [34]. The npy files can be obtained by running the script [34] on the original text file downloaded from [33]. The final dataset format used for the deep learning task thus consists of 905,684 data points333same as that used in the work [25], [28] of $16\times 20$ configuration matrices (whose entries are rescaled to be within the range $[0,1]$) and their associated Hodge number tuples ($h^{(1,1)}$,$h^{(2,1)}$, $h^{(3,1)}$, $h^{(2,2)}$). As noted in [25], there is a heavy imbalance in the dataset for $h^{2,1}$, as well as for $h^{3,1}$ and $h^{2,2}$: A large portion of the dataset for these Hodge numbers (on the order of $10^{5}$ data points) having values in the lower ranges. Furthermore, $h^{2,2}$ and $h^{3,1}$ (to a lesser extent, $h^{2,1}$) contain many outliers in the tail end of the data. These characteristics make the regression task for $h^{3,1}$, $h^{2,2}$ much harder than that for $h^{2,1}$ and $h^{1,1}$. Known results from the previous works [30], [25] have demonstrated the fact that although perfect accuracies could be obtained for $h^{1,1}$ and $h^{2,1}$, the accuracies for $h^{3,1}$ and $h^{2,2}$ are usually much lower.
@@
  • For the 5-fold CV training, the same datasets as the 72% data split were used, but with the following crucial differences:
    • The training set (72% of the data) is merged with the validation set (8% of the data), to form a total training set accounting for 80% of the data. This total training set is converted into 5 training datasets corresponding to the 5 CV training folds. Each ‘fold’ (where ’fold’ now is used synonymously with the training set for that fold) contains five partitions. In each of the training fold, a model is trained on the data of the 4 partitions (accounting for 64% of the full data) and validated on the remaning partition (accounting for 16% of the full data). For a 5-fold CV, there are 5 models. Although these models all have the same architecture, they should be treated as independent models, since they are trained and validated on different subsets of the training data.
@@
    • The distributions of the four Hodge numbers in each of the training folds are shown in Figs. 38, 39, 40, 41, 42 in Appendix 9.4. It can be seen from the figures there that the distributions in all five CV folds resemble the full dataset to the same degree, with none of the folds appearing significantly different from another.
@@ 3 Basics of recurrent neural networks
In this section, we briefly recall some pertinent mathematical facts about general recurrent neural networks (RNNs) and their more sophisticated GRU/LSTM variants for the benefits of the physicist readers who might not be familiar with deep learning. While our work makes use of both convolutional and recurrent architectures, we will not touch on the basics of convolutional neural networks, since many excellent texts already exist on CNN on the topic of deep learning CY manifolds as CNNs are the main tool for deep learning in these works (see, for example, the works [1], [5], [6], [4] which also discuss the basics of deep learning using a simple feed forward/dense articial neural network in great detail).
@@ 5 Training results (using 72% dataset)
In this section, we report the results obtained after the first round of training in which all twelve models was trained on the 72% dataset. Our training is done entirely on Kaggle (an AI/ML competition hosting platform) - https://www.kaggle.com - using the P100 GPU unit, which comes with a time usage limit888of 12 hours per notebook runtime session, (see https://www.kaggle.com/docs/notebooks) as well as a maximum of 30 hours per week, renewable at the start of each week. See https://www.kaggle.com/docs/efficient-gpu-usage. All networks were trained using HuberLoss [^9] https://pytorch.org/docs/stable/generated/torch.nn.HuberLoss.html, AdamW optimizer [^10] https://pytorch.org/docs/stable/generated/torch.optim.AdamW.html and a batch size of 128. The use of HuberLoss for training ensures a robust regression performance in the presence of outliers since HuberLoss combines the advantages of both the Mean Squared Error loss and the Mean Absolute Error (MAE) loss by using a $\delta$ parameter as follows:
@@ 5.1 CNN-GRU hybrid neural networks
  • For comparison among all CNN-GRU models, see Fig.16 which shows the validation accuracies obtained during training for all models in logarithmic scale. The training curves showing both training and validation losses for individual CNN-GRU models can be found in Fig.43.
    Figure 16: Validation losses of all CNN-GRU models (in log scale).
@@ 5.2 CNN-LSTM hybrid neural networks
  • The best performing CNN-LSTM hybrid model is CNN-LSTM-400 (with the train/test results noted in bold font). This is also the best model out of the 12 models considered in this work. A notable trend of these models’ results is the relatively stable performance observed at $M=256$ and $M=284$. Peak performance is attained at $M=400$. When increasing $M$ further to 416, the performance drops dramatically due to very slow convergence. In fact, the worst performing model of all models considered is CNN-LSTM-416, which also has the largest number of parameters ($3.017\times 10^{6}$). While all models achieved convergence after an average of around 300 epochs, CNN-LSTM-416’s training still progressed on at epoch 550 (the cutoff). If training was allowed to continue, the model would presumably continue to improve, but due to the time constraint with Kaggle’s GPU usage, it was not possible to extend the training time.
  • The obviously superior performance of LSTM-based hybrid networks over GRU-based ones can be seen when comparing the performances of CNN-LSTM-256 and CNN-LSTM-384 versus those of CNN-GRU-384 and CNN-GRU-416 (Fig.17). In particular, it is noted that the lighter CNN-LSTM-256 (at $1.55\times 10^{6}$ parameters) outperforms both the heavier CNN-GRU-384 ($2.22\times 10^{6}$) and CNN-GRU-416 ($2.49\times 10^{6}$). This is a clear indication of the efficiency of LSTM architecture over GRU’s in this task. In general, when comparing models with different underlying architectures, what matters most is the differences in the architectures rather than the number of parameters, since it is the architecture that determines how well a model can extract and learn from useful features [^13] While it might be fair to compare the performances of models from the same family (with the same underlying architectures) using solely their number of parameters, the same cannot be said for models from different families with different underlying architectures.. In this case, LSTMs with their more sophisticated double recurrent loops have proven better than their simpler GRU counterparts.
    Figure 17: Validation accuracies of CNN-LSTM-256 and CNN-LSTM-384 versus CNN-GRU-384/416 networks.
  • For comparison among the six CNN-LSTM models, see Fig.18 which shows the validation losses (log scale) of all CNN-LSTM models. The training curves showing training and validation losses for each individual CNN-LSTM model can be found in Fig.44 (in Appendix 9.5).
    Figure 18: Validation losses of all CNN-LSTM models (log scale). The lowest loss value (the green curve) corresponds to the best model CNN-LSTM-400. The red curve, corresponding to CNN-LSTM-416, did not drop to zero but ended at a very high value at the end of epoch 550 (signaling a failure to converge).
@@ 5.3 LSTM-based neural networks
  • For comparison among the four LSTM-based models, the validation losses of all four models are shown in Fig.19 in logarithmic scale.
    Figure 19: Validation losses of all LSTM models (log scale). In the plot, the green curve with the smallest loss corresponds to LSTM-448.
  • The training curves of individual LSTM-based models showing training and validation losses can be found in Fig.45.
@@ 5.4 Ensembles of best models
Forming ensemble is a common technique to improve the overall performance, especially when there are several trained models ready to be deployed. Due to the fact that each model has extracted their own useful features and thus learnt a different representation of the data, they might have different performance strengths in different parts of the data distribution. For instance, a model that performs well on a certain range of the data might not do well on a different range of the same data. In other words, each model has its own region of competence. In this sense, the diversity of model architectures in the ensemble potentially leads to a more robust performance and better generalization on the entire data set. In an ensemble, the predictions of all the constituent models are averaged, either evenly by taking the mean - as is the case in this section, or unevenly by taking an weighted average.
For the task at hand, the specific ensembles that we considered are listed in Table 5, together with their associated test accuracies (best results are listed in bold).